For a hydrogen atom, determine (a) the energy level corresponding to n 8; (b) whether there is an energy level at -2.500 x 10How do I solve this problem
Answer #1

Energy in hydrogen alom En = - 2.18 × 10-18 72 F9 = –2-18 X10 (8) Es = 3.40X10-20 J - 2.500 x 10-19 = -2.18 X 10 n2 n should